Scott Klasky

Office: NPAC, 3.183, Syracuse University, Syracuse , NY 13244-4100 Email: scott@npac.syr.edu; Phone: (315)-443-1690; Fax: (315)-443-1973

RESEARCH INTERESTS Numerical Relativity and applications to black hole spacetimes, numerical analysis with emphasis on multigrid techniques, numerical analysis tools for grand challenge applications. PRESENT POSITION PREVIOUS POSITIONS EDUCATION HONORS AND AWARDS RECENT RELEVANT PUBLICATIONS